If the NPV (net present value) of a shop is R195 000 and the profitability index is 1,2437, the initial investment is the is?


1
Expert's answer
2021-08-02T17:40:11-0400

"PI=\frac{(NPV+ i.investment)}{i.investment}"


"1.2437=\frac{(195 000+ i.investment}{( investment)}"

"1.2437=\frac{(195000+i.investment)}{initial investment}"


1.2437×i.investment=195000+i.investment

i.investment(1.2437-1)= 195000.

"0.2437i.investment = 195000"

i.investment = "\frac{195000}{0.2437}= 800164.136"